# RILe: Reinforced Imitation Learning

### v3

@@ 4 RILe: Reinforced Imitation Learning
To provide effective guidance, the discriminator must accurately identify whether a given state–action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$. GAIL [5] established the feasibility of such a discriminator (see Appendix B for details).
@@
$R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$ (7)
where $\upsilon(x)=2x-1$ scales the discriminator’s output, making it symmetric around zero. Including $a^{T}$ in the trainer’s reward ensures the trainer effectively learns from its own actions. Formally, the trainer’s objective is:
$\max_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
@@
RILe brings together these three components, student, trainer, and discriminator, to discover a student policy that imitates expert behaviors in $\tau_{E}$. Both $\pi_{S}$ and $\pi_{T}$ can be trained via any single-agent RL method. The overall training algorithm is detailed in Appendix J.
@@
$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[\upsilon(D_{\phi}(s^{T}_{t}))a^{T}_{t}]\right].$ (10)
By optimizing these objectives together, RILe efficiently learns both a reward function and a policy in high-dimensional settings where traditional AIL or IRL methods often struggle. Details on specific training strategies are provided in Appendix C.
@@ 5 Experiments
We evaluate the performance of RILe by addressing four key questions:
@@
    How does RILe’s reward-learning strategy differs from AIL and AIRL?
@@
    How adaptive is RILe’s learned reward function?
@@
    How does RILe perform in both low- and high-dimensional continuous-control problem compared to AIL/AIRL/IRL?
@@
    How explicitly using expert-data within RILe’s training process affects the context-sensitive reward learning?
Baselines We compare RILe with seven baseline methods: Behavioral cloning (BC [4, 21], BCO [10]), adversarial imitation learning (GAIL [5], GAIfO [9] and DRAIL [11]), adversarial inverse reinforcement learning (AIRL [8]), and inverse reinforcement learning (IQ-Learn [7]). DRAIL [11] introduces a diffusion-based discriminator implementation, which is applied to both GAIL and RILe, and referred as DRAIL-GAIL and DRAIL-RILe.
Additional experimental details are provided in the Appendix D, and hyperparameter selections are discussed in the Appendix H.
5.1 Evolving Reward Function
To answer the first question about RILe’s reward-learning strategy, we compare RILe’s performance with AI(R)L baselines in a maze setting. In this environment, the agent must navigate from a fixed start to a goal while avoiding static obstacles; we use a single expert demonstration.
@@ 5.1 Evolving Reward Function
Figure 3: Reward Function Comparison. Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment. Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right). The expert’s trajectory is shown in red, while the student agent’s trajectory from the previous training epoch is in black. Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards. Purple squares represent obstacles. RILe demonstrates a dynamic reward function that adapts with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
@@
RILe’s reward function dynamically adapts to the student’s current policy, providing guidance that encourage suboptimal actions which eventually lead the student closer to the expert trajectory. By contrast, GAIL and AIRL’s reward functions remain relatively static. Specifically, the first column in Figure 3 shows RILe’s trainer encouraging exploration toward the bottom-right of the maze, which is initially suboptimal but helpful in the long run. As the student learns to reach the lower part of the maze, RILe shifts high-reward regions toward the top-left (second column), again encouraging incremental progress. The third column illustrates how RILe boosts rewards near the goal while still maintaining some incentive around other key areas to keep the agent from getting stuck. Overall, RILe’s evolving reward function serves as a curriculum that promotes gradual improvement toward expert-like performance. This dynamic reward adaptation gets important in higher-dimensional tasks as we show in Section 5.3.
5.2 Reward Function Dynamics
To address the second question on adaptability, we quantify how the reward function evolves and its relationship to the student’s performance. We compare RILe with GAIL, DRAIL-GAIL, and DRAIL-RILe in a high-dimensional robotic control scenario (learning to walk with UnitreeH1 robot).
@@ 5.2 Reward Function Dynamics
5.2.1 Adaptability of the Learned Reward Function
We assess how adaptive the reward function learned by RILe is compared to that of AIL. Fig. 4(a) presents changes in reward distributions over 10,000 consecutive steps. RILe exhibits the highest adaptability in its reward function, aligning with our goal of having the reward function adapt based on the student’s learning stage. The advanced discriminator in DRAIL reduces the need for drastic reward function changes, yet RILe remains more adaptive than GAIL. Since the changing student policy indirectly affects the Fig. 4(a), we also show changes in reward values for the fixed set of states in Fig. 4(b). Again, RILe’s reward function is the most adaptive among all methods.
Higher adaptability of RILe ensures that the reward signal remains aligned with incremental performance improvements, enabling the student to receive more timely and effective guidance throughout training.
5.2.2 Correlation between the Learned Reward and the Student Performance
@@
5.3 Imitation in Continuous Control Tasks
5.3.1 Motion-Capture Data Imitation for Robotic Continuous Control
To answer the third question, we first test RILe in high-dimensional robotic locomotion tasks [22], where the agent must imitate motion-capture data for various robotic bodies. This benchmark is especially demanding due to its complexity and dimensionality.
@@ 5.3.1 Motion-Capture Data Imitation for Robotic Continuous Control
Table 1 presents results for seven LocoMujoco tasks across different test seeds (see Appendix D.3 for details). Overall, these results underscore that RILe outperforms the AIL/IL/IRL baselines, and benefits even more from the enhancements provided by the DRAIL variants, achieving performance levels close to the expert. The performance variations across tasks indicate that although RILe’s adaptive reward function is highly effective, task-specific factors also play a role, suggesting avenues for future research. This overall performance aligns with our claim that an adaptive reward function is crucial for mastering complex, high-dimensional behaviors
Table 2: Test results on four MuJoCo tasks.
RILe GAIL AIRL IQ
Humanoid 5928 5709 5623 327
Walker2d 4435 4906 4823 270
Hopper 3417 3361 3014 310
HalfCheetah 5205 4173 3991 755
Figure 5: Explicit Usage of Expert Data. Red and yellow markers show normalized scores and steps, respectively. Expert data usage speeds the training of RILe but reduces final performance.
5.3.2 Learning from Demonstrations
We further evaluate RILe’s performance on four standard MuJoCo tasks [23, 24], Humanoid, Walker2d, Hopper, and HalfCheetah, with baseline comparisons widely reported in prior work.
Table 2 summarizes test results. Overall, these results demonstrate that RILe consistently achieves competitive or superior performance compared to established AIL/AIRL/IRL methods. In particular, on high-dimensional tasks such as Humanoid, RILe’s performance advantage is evident, underscoring the effectiveness of its adaptive reward function in handling complex continuous control challenges.
5.4 Explicit Usage of Expert Data
To answer the fourth question, we use the Humanoid environment in MuJoCo [23, 24] with a single expert trajectory from [7]. We vary the proportion of expert data in the replay buffers from 0% to 100%. For instance, 25% indicates a quarter of each buffer is expert data and 75% is agent data (see Appendix D.5 for details).
Figure 5 illustrates how changing the proportion of expert data in both the trainer’s and the student’s replay buffers influences RILe’s training. Introducing more expert data accelerates RILe’s convergence (needs fewer training steps), but reduce the final performance. In the extreme case of 100% expert data, the student’s performance drops substantially. This indicates that too much expert data hamper the trainer’s ability to adapt to the student’s real-time needs, disrupting RILe’s context-sensitive reward customization.
We also compare RILe’s results with IQ-Learn and BC, both of which rely heavily on expert data. Even when RILe is trained with substantial amounts of expert data, it still generally performs better, indicating that RILe’s adaptive reward-shaping provides a crucial edge over supervised methods.
@@ 6 Discussion
Our maze experiments provide an interpretable example of how the trainer agent tailors its rewards. By encouraging actions that might seem suboptimal for immediate imitation but advantageous for long-term learning, RILe establishes a curriculum that ultimately boosts performance. This adaptive strategy helps RILe to achieve superior results in our continuous control experiments, where reward shaping becomes especially critical in high-dimensional settings.
Nonetheless, policy stability remains challenging when the reward function is constantly evolving. Freezing the trainer (see Appendix C) stabilizes learning but halts further adaptation, and the discriminator itself tends to overfit quickly. Future work could explore fully cooperative multi-agent RL techniques to allow ongoing adaptation, investigate ways to bound or regulate trainer updates, and consider discriminator-less formulations for reward learning.
@@ B.1 Lemma 1:
The trainer aims to maximize $Q_{T}(s^{T}_{t},a^{T}_{t})$ to satisfy Equation 13. Since $r^{T}_{t}$ depends directly on $D_{\phi}(s^{T}_{t})$ and $a^{T}_{t}$, the trainer learns to select $a^{T}_{t}$ that maximizes $Q_{T}(s^{T}_{t},a^{T}_{t})$ Considering that $a^{T}_{t}\in[-1,1]$, the immediate reward $r^{T}_{t}$ is maximized when $a^{T}_{t}$ has the same sign as $\upsilon(D_{\phi}(s^{T}_{t}))$ Therefore, the optimal action $a^{*T}_{t}$ is:
$\alpha^{*T}_{t}=\begin{cases}1,&\text{if }D\phi(s_{t}^{T})>0.5,\\ -1,&\text{if }D\phi(s_{t}^{T})<0.5,\\ \text{any value in }[-1,1],&\text{if }D\phi(s_{t}^{T})=0.5.\end{cases}$ (16)
Equation 16 implies the trainer assigns positive rewards to student state-action pairs that the discriminator assesses as more likely to be from the expert ($D_{\phi}(s^{T}_{t})>0.5$) and negative rewards to those unlikely to be from the expert ($D_{\phi}(s^{T}_{t})<0.5$). By this mechanism, the trainer’s policy optimization relies on the discriminator’s assessment to assign rewards that encourage expert-like behavior.
@@
$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})\left((2D_{\phi}(s_{t}^{T})-1)a_{t}^{T}+\gamma Q_{T}(s_{t+1}^{T},a_{t+1}^{T})\right)\right]$ (17)
The trainer adjusts $\pi_{T}$ to output high rewards when $D_{\phi}(s_{t}^{T})$ is high. Therefore the trainer learns to assign higher rewards to student behaviors that are more similar to the expert behaviors, according to the discriminator.

### v4

@@ 4 RILe: Reinforced Imitation Learning
To provide effective guidance, the discriminator must accurately identify whether a given state–action pair originates from the expert distribution $(s,a)\sim\tau_{E}$ or not $(s,a)\notin\tau_{E}$. GAIL (Ho & Ermon, 2016) established the feasibility of such a discriminator (see Appendix B for details).
@@
$R^{T}=e^{|\upsilon(D_{\phi}(s^{T}))-a^{T}|}$ (7)
where $\upsilon(x)=2x-1$ scales the discriminator’s output, making it symmetric around zero. Including $a^{T}$ in the trainer’s reward ensures the trainer effectively learns from its own actions. Formally, we define the trainer’s objective as:
$\max_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[e^{|\upsilon(D_{\phi}(s^{T}))-a^{T}|}].$ (8)
@@
RILe brings together these three components, student, trainer, and discriminator, to discover a student policy that imitates expert behaviors in $\tau_{E}$. Both $\pi_{S}$ and $\pi_{T}$ can be trained via any single-agent RL method. The overall training algorithm is detailed in Appendix I.
@@
$\pi_{T}^{*}=\operatorname*{arg\,max}_{\pi_{T}}\mathbb{E}_{\begin{subarray}{c}s^{T}\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}[e^{|\upsilon(D_{\phi}(s_{t}^{T}))-a_{t}^{T}|}]\right].$ (10)
By optimizing these objectives together, RILe efficiently learns both a reward function and a policy in high-dimensional settings where traditional AIL or IRL methods often struggle. Achieving stable joint training requires specific techniques in RILe’s adaptive system, which are discussed in Appendix C.
@@ 5 Experiments
We evaluate the performance of RILe in this section. We perform five ablation studies to analyze RILe:
@@
    Reward Function Evaluation: We qualitatively analyze how RILe’s reward-learning strategy differs from baselines.
@@
    Reward Function Dynamics: We quantitatively assess the reward function learned by RILe.
@@
    Trainer-Discriminator Relation: We compare different trainer reward functions and investigate how the interaction between the trainer and the discriminator affects RILe’s performance
@@
    Robustness to Noise and Covariance Shift: We investigate the robustness of RILe to different types of noise and covariate shift in the environment.
  5. 5.
    Explicit Usage of Expert Data: Analyzing the effect of using expert data explicitly inside RILe on the trainer-student dynamics.
Then, we evaluate the performance of RILe in high-dimensional tasks with two experiments:
  1. 1.
    Learning from Expert Demonstrations: Learning from perfect expert demonstrations in a reinforcement learning benchmark.
  2. 2.
    Robotic Continuous Control from Motion Capture Data: Learning to walk and to carry a box from motion-capture data with four different embodiments with unique configurations.
@@
Figure 3: Reward Function Comparison. Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment. Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right). The expert’s trajectory is shown in red, while the student agent’s trajectory from the previous training epoch is in black. Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards. Black squares represent obstacles. RILe demonstrates a dynamic reward function that adapts with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training and struggle to adapt.
Baselines We compare RILe with seven baseline methods: Behavioral cloning (BC (Bain & Sammut, 1995; Ross & Bagnell, 2010), BCO (Torabi et al., 2018a)), adversarial imitation learning (GAIL (Ho & Ermon, 2016), GAIfO (Torabi et al., 2018b) and DRAIL (Lai et al., 2024)), adversarial inverse reinforcement learning (AIRL (Fu et al., 2018)), and inverse reinforcement learning (IQ-Learn (Garg et al., 2021)). DRAIL (Lai et al., 2024) introduces a diffusion-based discriminator implementation, which is applied to both GAIL and RILe, and referred as DRAIL-GAIL and DRAIL-RILe.
Additional experimental details are provided in the Appendix D, and hyperparameter selections are discussed in the Appendix G.
5.1 Ablation Studies
5.1.1 Reward Function Evaluation
To qualitatively evaluate how RILe’s reward-learning strategy differs from AI(R)L baselines, we compare them in a maze environment. In this environment, the agent must navigate from a fixed start to a goal while avoiding static obstacles; we use a single expert demonstration.
@@ 5.1.1 Reward Function Evaluation
RILe’s reward function dynamically adapts to the student’s current policy, providing guidance that encourage suboptimal actions which eventually lead the student closer to the expert trajectory. By contrast, GAIL and AIRL’s reward functions remain relatively static. Specifically, the first column in Figure 3 shows RILe’s trainer encouraging exploration toward the bottom-right of the maze, which is initially suboptimal but helpful in the long run. As the student learns to reach the lower part of the maze, RILe shifts high-reward regions toward the top-left (second column), again encouraging incremental progress. The third column illustrates how RILe boosts rewards near the goal while still maintaining some incentive around top-left areas to keep the agent from getting stuck.
Overall, RILe’s evolving reward function serves as a curriculum that promotes gradual improvement toward expert-like performance by encouraging exploration. This dynamic reward adaptation gets important in higher-dimensional tasks as we show in Section 5.3.
@@
5.1.2 Reward Function Dynamics
To qualitatively evaluate how RILe’s reward function evolves during the training, we compare RILe with GAIL, DRAIL-GAIL, and DRAIL-RILe in a high-dimensional robotic control scenario (learning to walk with UnitreeH1 robot).
Adaptability of the Learned Reward Function
We compare the adaptability of the reward functions learned by RILe and AIL. Fig. 4(a) presents changes in reward distributions over 10,000 consecutive steps. RILe exhibits the highest adaptability in its reward function, aligning with our goal of having the reward function adapt based on the student’s learning stage. The advanced discriminator in DRAIL reduces the need for drastic reward function changes, yet RILe remains more adaptive than GAIL. Since the changing student policy indirectly affects RFDC, we also show changes in reward values for the fixed set of states in Fig. 4(b). Again, RILe’s reward function is the most adaptive among all methods. Higher adaptability of RILe ensures that the reward signal remains aligned with incremental performance improvements, enabling the student to receive more timely and effective guidance throughout training.
Correlation between the Learned Reward and the Student Performance
@@ 5.1.2 Reward Function Dynamics
Figure 5: Trainer-Discriminator Relation: Comparison of different trainer reward functions, each defining a different relationship between the trainer’s action and the discriminator’s output. The student’s return curves on the left show how performance evolves, and the normalized final performance on the right presents a clear comparison between reward designs. Exponential naive converges faster but plateaus at a lower final reward, whereas exponential difference yields the highest performance.
5.1.3 Trainer-Discriminator Relation
We investigate how the interaction between the trainer agent and the discriminator affects RILe’s performance by comparing different trainer reward functions. Each reward function defines a different relationship between the trainer’s action $a^{T}$ and the discriminator’s output $D_{\phi}(s^{T})$. We consider following reward functions: (a) Difference ($R^{T}=-|\upsilon(D_{\phi}(s^{T}))-a^{T}|$), (b) Exponential Difference (default in RILe): $R^{T}=e^{-|\upsilon(D_{\phi}(s^{T}))-a^{T}|}$, (c) Multiplication ($R^{T}=\upsilon(D_{\phi}(s^{T}))a^{T}$)), (d) Naive ($R^{T}=D_{\phi}(s^{T})$)), (e) Exponential Naive ($R^{T}=e^{1-D_{\phi}(s^{T})}$)) and (f) Sigmoid (($R^{T}=D_{\phi}(s^{T})\sigma(a^{T})$))), where $\upsilon(x)=2x-1$ and $\sigma(x)=\frac{1}{1+e^{-x}}$.
Figure 5 presents reward curves and normalized rewards, where all rewards are normalized according to the maximum mean achieved test reward. While the exponential naive reward function offers the fastest convergence, the exponential difference reward offers the best performance. Therefore, we use exponential difference reward as the default reward function in RILe.
5.1.4 Robustness to Noise and Covariate Shift
We evaluate RILe’s robustness to noisy expert demonstrations and environmental covariate shift. First, to evaluate the noise robustness of RILe In the MuJoCo Humanoid-v2 environment, we inject a zero-mean Gaussian noise (varying $\Sigma$) into either expert actions or states. We use GAIL (Ho & Ermon, 2016), AIRL (Fu et al., 2018), RIL-Co (Tangkaratt et al., 2021), IC-GAIL (Wu et al., 2019), and IQ-Learn (Garg et al., 2021) as baselines. Table 2 shows that RILe consistently outperforms baselines across all noise levels. Notably, it maintains high performance even under heavy noise ($\Sigma=0.5$).
Second, inspired by Xu et al. (2022), we evaluate the stability of the reward functions learned by RILe and AIRL. First, we train both models in a clean environment. Then, we freeze the learned reward functions and train new student agents in environments where Gaussian noise is injected into their actions (covariate shift). Table 2 shows that the reward function learned by RILe demonstrates superior robustness to covariate shift, maintaining high performance even under increased noise levels.
Table 1: Robustness to different noise levels in the expert demonstrations in MuJoCo Humanoid-v2 environment.
RILe GAIL AIRL RIL-Co IC-GAIL IQ
Noise Free $\Sigma=0$ 5928 5709 5623 576 610 327
Action $\Sigma=0.2$ 5280 5275 4869 491 601 192
$\Sigma=0.5$ 5154 902 4589 493 568 153
State $\Sigma=0.2$ 5350 5147 4898 505 590 243
$\Sigma=0.5$ 5205 917 4780 501 591 277
Table 2: Robustness to covariate shifts in environment.
RILe AIRL
No Noise 5928 5623
Mild
$\Sigma=0.2$ 5201 5005
High
$\Sigma=0.5$ 5196 4967
5.1.5 Explicit Usage of Expert Data inside RILe
Figure 6: Explicit Usage of Expert Data. Red and yellow markers show normalized scores and steps. Expert data usage speeds the training of RILe but reduce final performance.
To assess the impact of using expert data explicitly inside RILe, we experiment in the MuJoCo Humanoid environment using a single expert trajectory from (Garg et al., 2021). We vary the proportion of expert data in the replay buffers from 0% to 100% (e.g., 25% indicates one quarter of each buffer is expert data; see Appendix D.5).
Figure 6 shows that while more expert data in both the trainer’s and the student’s replay buffers accelerates RILe’s convergence, it reduces final performance. At 100% expert data, the student’s performance drops markedly. This indicates that excessive expert data hampers the trainer’s real-time adaptation, disrupting RILe’s context-sensitive reward customization.
We compare RILe with IQ-Learn and BC, both of which rely heavily on expert data. Even with substantial amounts of expert data, RILe still performs better than baselines, indicating that RILe’s adaptive reward-shaping provides a crucial edge over those methods.
Table 3: Test results on four MuJoCo tasks.
RILe GAIL AIRL IQ
Humanoid 5928 5709 5623 327
Walker2d 4435 4906 4823 270
Hopper 3417 3361 3014 310
HalfCheetah 5205 4173 3991 755
5.2 Learning from Expert Demonstrations
We evaluate how RILe’s performance compares to baselines in a widely used learning from demonstration benchmark (Todorov et al., 2012; Brockman et al., 2016), where perfect expert state-action pairs are available.
Table 3 shows RILe consistently achieves competitive or superior performance compared to AIL/AIRL/IRL methods. In particular, on high-dimensional tasks (e.g. Humanoid), RILe’s performance advantage is evident, underscoring the effectiveness of its adaptive reward function.
5.3 Robotic Continuous Control from Motion Capture Data
We test RILe in high-dimensional robotic locomotion tasks (Al-Hafez et al., 2023),for various robotic bodies. The agent must imitate motion-capture data, which is inherently noisy and only consists states without action information. This benchmark is especially demanding due to its complexity and dimensionality.
@@ 5.3 Robotic Continuous Control from Motion Capture Data
Table 4 presents results for seven LocoMujoco tasks across different test seeds (see Appendix D.3 for details). Overall, these results underscore that RILe outperforms the AIL/IL/IRL baselines, and benefits even more from the enhancements provided by the DRAIL variants, achieving performance levels close to the expert. The performance variations across tasks indicate that although RILe’s adaptive reward function is highly effective, task-specific factors also play a role. This overall performance aligns with our claim that an adaptive reward function is crucial for mastering complex, high-dimensional behaviors
@@ 6 Discussion
Our maze experiment illustrates how the trainer agent tailors its rewards. By encouraging actions that might seem suboptimal for immediate imitation but advantageous for long-term learning, RILe establishes a curriculum that ultimately boosts performance. This adaptive strategy helps RILe achieve superior results in our continuous control experiments, where reward shaping becomes especially critical in high-dimensional settings.
Nonetheless, policy stability remains challenging with dynamically evolving rewards. Freezing the trainer (see Appendix C) stabilizes learning but halts further adaptation, and the discriminator itself tends to overfit quickly. Future work may explore cooperative multi-agent RL to support continual adaptation, and consider discriminator-less formulations for reward learning.
@@ B.1 Lemma 1:
The trainer aims to maximize $Q_{T}(s^{T}_{t},a^{T}_{t})$ to satisfy Equation 13. Since $r^{T}_{t}$ depends directly on $D_{\phi}(s^{T}_{t})$ and $a^{T}_{t}$, the trainer learns to select $a^{T}_{t}$ that maximizes $Q_{T}(s^{T}_{t},a^{T}_{t})$ Considering that $a^{T}_{t}\in[-1,1]$, the immediate reward $r^{T}_{t}$ is maximized when $a^{T}_{t}$ matches $\upsilon(D_{\phi}(s^{T}_{t}))$ Therefore, the optimal action $a^{*T}_{t}$ is:
$\alpha^{*T}_{t}=\upsilon(D_{\phi}(s^{T}_{t}))=2D_{\phi}(s^{T}_{t})-1$ (16)
Equation 16 implies the trainer learns to match the discriminator’s scaled output. By this mechanism, the trainer’s policy optimization relies on the discriminator’s assessment to assign rewards that encourage expert-like behavior. Over time, this guides the student toward regions of the state-action space where $D_{\phi}(s^{T}_{t})\approx 1$, i.e., expert-like behavior.
@@
$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})\left(e^{-|\upsilon(D_{\phi}(s_{t}^{T}))-a_{t}^{T}|}+\gamma Q_{T}(s_{t+1}^{T},a_{t+1}^{T})\right)\right]$ (17)
The trainer adjusts $\pi_{T}$ to output high rewards when $D_{\phi}(s_{t}^{T})$ is high. Therefore the trainer learns to assign higher rewards to student behaviors that are more similar to expert behaviors, according to the discriminator.